In the hydrolysis of triacylglycerol esters (TAE) known as saponification.

1 mole of triglyceride requires 3 moles of NaOH to produce 1 mole of glycerol and 3 molecule of soap ( Carboxylate salt).

Start with Mass of vegetable oil: 5.007 g

Mass of NaOH : 5.008g

Question:1. Determine the limiting reactant. Show the work

2. What is the percent yield of product? ( Mass of the soap recovered= 10.95g)
Answer #1

The reaction (balanced ) is:

1 ( triglyceride) + 3 NaOH \rightarrow 1( glycerol) + 3 (soap)

C57H104O6 + 3 NaOH \rightarrow    C3H8O3 + 3 C18H33NaO2

Molar mass of triglyceride, that is vegetable oil ( C57H104O6 ) = 885.45 g/mol  

So, moles of triglyceride present = ( 5.007 g ) / (885.45 g /mol) = 0.005655

For this many moles, 3 times mole of NaOH are required; = 3 x 0.005655 = 0.01696

Molar mass of NaOH = 40.00 g/mol

So, moles of NaOH present = ( 5.008 g ) / (40.00 g /mol) = 0.1252

Excess moles of NaOH are present and excess = 0.1252 - 0.01696 = 0.1082

Therefore triglyceride is the limiting reagent.

The mole of product would be determined as per the mole of triglyceride present. As per the reaction stoichiometry, 1 mole of triglyceride gives 3 mole of soap,

Therefore 0.005655 mole of triglyceride present would give 3 x 0.005655 = 0.01696 mole of soap.

To determine the molar mass of soap obtained from the vegetable oil;

Molar mass of soap (C18H33NaO2 )=   304 g/mol

Amount of soap produced = (0.01696 mole) x 304 g/mol    = 5.156 g (theroretical yield)

This is problematic, as theroretical yield should be greater than experimental yield (= 10.95g ). That means, there is some issue with the data. Or, mention the exact type of vegetable oil used, which would change the number of moles calculation, and answer will be different.

The method of calculation would remain same as described above.

The % yield can then be calculated as : { (theroretical yield - experimental yield ) / theroretical yield } x 100